Determine overall cell potential for a system in which the following equation represent the net reaction zno (s)+fe++<>zn+++feo (s)
1
Expert's answer
2019-08-19T02:06:42-0400

In here Zn(II) is getting converted to Zn.

and Fe is getting converted to Fe(II)

Thus (1) Zn2+ + 2e- = Zn , E= -0.76 V

(2) Fe = Fe (II) + 2e- , E= 0.44

(1) + (2) = E= -0.76+0.44 = -0.22 V